OpenStudy (anonymous):

solve for x: x + sqrt(x) - 12 = 0

OpenStudy (anonymous):

x= 3is you answer.

OpenStudy (anonymous):

firstly let x = [y ^{2}\]

OpenStudy (anonymous):

and then solve the new form of equation : \[y ^{2}+y-12 = 0\] and find the answers for y after that omit the unsuitable answers for x those which are negative I mean.

OpenStudy (anonymous):

was it useful ?

OpenStudy (anonymous):

\[y ^{2}+y -12 = 0 \] => \[(y+4)(y-3)= 0\] so y can be 3 or -4.but x cannot be sqrt(-4) so x is sqrt(3).

OpenStudy (lgbasallote):

actually @mas_gh90 ... if y = 3 and y = -4...then sq x = 3 and sq x = -4 so x = 9 and x =16

OpenStudy (anonymous):

that doesn't make sense, can we just make x = y^2?

OpenStudy (lgbasallote):

yes...x = y^2..so y = sqrt of x...since the final answers involve y and not y^2 you substitute back to sqrt x then square it to get x. Do you get it? :D

OpenStudy (anonymous):

you're completely right dear lgbasallote,sorry for my foolish mistake.

OpenStudy (anonymous):

alright i get it thank you both!

OpenStudy (anonymous):

Don't mention it =)

OpenStudy (anonymous):

x +√x - 12 = 0 √x = 12 -x ->x = ( 12 - x)² x = 144 - 24x + x² x² - 25x + 144 = 0 => x = 16, x = 9